A chi-square independence test is to be conducted to decide whether an association exists between two variables of a populations. One variable has 4 possible values and the other variable has 6. What are the degrees of freedom for the 𝜒2 -statistic?

answer:


1
Expert's answer
2021-12-09T10:17:28-0500

Let the variable with 4 possible values be on the columns side and the variable with 6 possible values be on the rows side. We need to determine the number of rows and the number of columns as follows.

Number of columns=4

Number of rows =6

The formula to determine the number of degrees of freedom is, "df=(r-1)*(c-1)", where "r" is the number of rows and "c" is the number of columns. Therefore, the value of degrees of freedom is, "df=(6-1)*(4-1)=5*3=15"

Thus degrees of freedom, df=15.